OpenStudy (fools101):

Find the simplest polynomial function with the given zeros. + or - 2, 3i A) P(x) = x4 - 4x3 + 9x2 - 36 B) P(x) = x4 - 13x2 - 36 C) P(x) = x4 + 5x2 - 36 D) none of the above

OpenStudy (fools101):

please help!

OpenStudy (fools101):

I think is C!!!!?

OpenStudy (anonymous):

Well, it's gonna be (x - 2)(x + 2)(x - 3i)(x + 3i). So if you multiply that out, you get: \[(x^2 - 4)(x^2 + 9)\] = \[x^4 + 9x^2 - 4x^2 - 36\] = \[x^2 + 5x^2 - 36\] So yeah, you got it. Good job! =)

OpenStudy (anonymous):

Err...I mean x^4 on the last one =\

OpenStudy (fools101):

So it D? HUH???

OpenStudy (fools101):

????

OpenStudy (anonymous):

I'm sorry, I meant \[x^4 + 5x^2 - 36\] So it's C. I mis-typed.

OpenStudy (anonymous):

C.\[x^4+5 x^2-36=(x-2) (x+2) \left(x^2+9\right) \]

OpenStudy (fools101):

AWWWW SWEET IT OK THSNK U TNSGOD thank u for helping me make sure it was right!!! THANK U ROBTOBEY : D!!!!

OpenStudy (fools101):

Thank u ***

OpenStudy (anonymous):

Your are welcome.

OpenStudy (fools101):

^ ^..!!!